OpenStudy (anonymous):

How do I solve for x? (2-x)/6 - (x/3) = 1 - (2x-2)/2

OpenStudy (gw2011):

(2-x)/6-(x/3)=1-[(2x-2)/2] Multiply the equation by 6 to eliminate the denominators 2-x-2x=6-3(2x-2) -3x+2=6-6x+6 -3x+6x=6+6-2 3x=10 x=10/3

OpenStudy (anonymous):

x=10/3